An electron is traveling at 1.9x10 m/s when it enters an electric field that slows it. The field has a mag nitude of 6200 N/C. How far does the electron penetrate the field before it stops.


1
Expert's answer
2022-04-21T13:01:45-0400

The energy-work theorem says

"\Delta E=W\\
\frac{mv^2}{2}=Fd"

Hence

"d=\frac{mv^2}{2F}=\frac{9.1*10^{-31}*(1.9*10)^2}{2*6200}=2.6*10^{-32}\;\rm m"